OpenStudy (anonymous):

Melinda makes monthly car payments of $285 each month. How much will she have paid after 2 years? 5 years?

OpenStudy (anonymous):

see 1 month = 1/12 year

OpenStudy (anonymous):

therefore ratios might help... 285 is paid in 1/12 year how much will be paid in 2 years... 285:1/12 x : 2 now cross multiply.. 1/12*x=285*2 x=570*12 or x=570/(1/12) x= $6840

OpenStudy (anonymous):

now do the same will 5 years... 285:1/12 x : 5 now cross multiply.. 1/12*x=285*5 x=1425*12 or x=1425/(1/12) x= $17100

OpenStudy (anonymous):

i hope it helped!